OpenStudy (anonymous):

use the given slope and point to graph each line. 1 - (2,1) 2,

OpenStudy (anonymous):

y-1=1/2(x-2) y=1/2x-1+1 = 1/2x y=1/2x

OpenStudy (anonymous):

y - 1 = 1/2(x - 2) y = 1/2x - 1 + 1 y = 1/2x

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

No problem.